If there are 6500 degree-days in this area, and the price of fuel is $10 per...

  1. If there are 6500 degree-days in this area, and the price of fuel is $10 per million Btu, then find the heating cost per square foot per heating season for the un-insulated floor. (ans: $0.68/ft2)
  2. If 6-in fiberglass costs $0.40 per square foot, what will be the payback time (as a result of energy savings) on this installation? (ans: 0.7 seasons)
Answer #1

a)

24hrs / dau × D D 1ーー× 24 × 6500 67826.1 Btu/ft annual- A R

Heating Cost

10 67826.1 × ĩ06

=$0.678/ft2 =$ 0.68 /ft2

b)

Q 1 = -x A R 24hrs/day × D Dannual × 24 × 6500 = 8210.5263 Btu/ft2

Savings per season=0.68-8210.5263*(10/106)=$ 0.596 /season

Payback time =0.4/0.596 =0.671 seasons =0.7 seasons (approx)